# Eka-Eval : A Comprehensive Framework for Indian Large Language Models

### v3

@@ Abstract
The rapid advancement of Large Language Models (LLMs) has intensified the need for evaluation frameworks that address the requirements of linguistically diverse regions, such as India, and go beyond English-centric benchmarks. We introduce Eka-Eval, a unified evaluation framework that integrates over 35+ benchmarks (including 10 Indic benchmarks) across nine major evaluation categories. The Framework provides broader coverage than existing Indian language evaluation tools, offering 11 core capabilities through a modular architecture, seamless integration with Hugging Face and proprietary models, and plug-and-play usability. As the first end-to-end suite for scalable, multilingual LLM benchmarking, the framework combines extensive benchmarks, modular workflows, and dedicated support for low-resource Indian languages to enable inclusive assessment of LLM capabilities across diverse domains. We conducted extensive comparisons against five existing baselines, demonstrating that Eka-Eval achieves the highest participant ratings in four out of five categories.
[…]
Eka-Eval : A Comprehensive Framework for Indian Large Language Models
@@ 1 Introduction
Large Language Models (LLMs) have rapidly transformed natural language processing (NLP), enabling impressive generalization across diverse tasks including instruction following, reasoning, summarization, translation, and tool use. With the advent of general-purpose foundation models such as GPT-4 (Achiam et al., 2023), Claude (Anthropic, 2024), Gemini (Anil et al., 2023) and Llama-3 (Touvron et al., 2023), the focus of research has increasingly shifted from building task-specific models to systematically evaluating these powerful systems. Evaluation plays a critical role not only in measuring progress but in identifying capabilities, exposing limitations, and informing deployment strategies.
@@
In response, several evaluation frameworks have emerged, including HELM (Liang et al., 2022), EleutherAI Harness (Gao et al., 2021), lm-eval-harness (Gao et al., 2021), and OpenCompass (OpenCompass Contributors, 2023). However, existing evaluation frameworks are primarily centered on English or other high-resource languages and offer limited support for low-resource or multilingual settings (Watts et al., 2024). This lack of inclusivity significantly limits the effectiveness of these tools in linguistically diverse regions like India, which has 22 constitutionally recognized languages encompassing over a billion native speakers.
While notable benchmarks like IndicGLUE (Kakwani et al., 2020), IndicGenBench (Singh et al., 2024), and MILU (Verma et al., 2024) have addressed some gaps in Indic-language evaluation, they remain siloed efforts and lack integration into unified evaluation workflows — requiring users to perform manual configuration to use them in real-world LLM assessment pipelines.
Furthermore, popular frameworks such as FreeEval (Yu et al., 2024) and lm-eval-harness (Gao et al., 2021) require extensive configuration and engineering expertise, limiting their adoption by developers and researchers operating in low-resource environments. These challenges create a need for an actively maintained, community driven, multilingual, and task-diverse evaluation suite.
Framework Custom Datasets Custom Models Custom Prompting Long Context Tool Use Distributed Inference Visual Analysis Multilingual Indic Production Optimization Interactive CLI Quantization
@@
Table 1: Comparison of LLM Evaluation Frameworks across eleven key capabilities, showing the comprehensive coverage of Eka-Eval.
To address these limitations, we introduce Eka-Eval, a unified, extensible, and ready-to-use evaluation framework for LLMs, integrating over 35 benchmarks spanning both global (English) and Indic-language tasks. The overall architecture, presented in Figure 1, is designed to be modular, compatible with Hugging Face and proprietary models, it offers plug-and-play usability with minimal configuration overhead. To contextualize Eka-Eval’s capabilities, Table 1 compares leading LLM evaluation frameworks across eleven key capabilities, demonstrating its more comprehensive feature coverage relative to existing frameworks.
Eka-Eval covers nine major evaluation categories: (i) Code Generation and programming; (ii) Mathematics and logical reasoning; (iii) Reading comprehension; (iv) Commonsense reasoning; (v) World knowledge; (vi) Long-context understanding; (vii) General reasoning and knowledge; (viii) Tool use and API reasoning; and (viiii) Indic-specific NLP benchmarks including ARC-C (Clark et al., 2018), BoolQ (Clark et al., 2019), GSM8K (Cobbe et al., 2021), Math (Hendrycks et al., 2021), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2020a), SQuAD (Rajpurkar et al., 2018), APIBench (Patil et al., 2024), IndicGenBench (Singh et al., 2024).
Our key contributions are:
@@
  • We propose Eka-Eval, a unified and modular evaluation framework that integrates over 35 benchmarks across nine major evaluation categories, providing comprehensive assessment of both global and Indic-language LLM capabilities.
  • We provide comprehensive support for most Indian languages and conduct thorough human evaluation across five existing frameworks on four benchmark tasks, demonstrating Eka-Eval’s effectiveness and reliability.
2 Related Work
The evaluation of LLMs has evolved from task-specific benchmarks to comprehensive, modular frameworks that support diverse capabilities and deployment settings. This section categorizes prior work into three major areas: general-purpose evaluation frameworks, specialized capability benchmarks, and multilingual/Indic evaluations. We conclude by highlighting how Eka-Eval integrates and advances these directions.
General-Purpose LLM Evaluation Frameworks: Early benchmark suites such as GLUE (Wang et al., 2018) and SuperGLUE (Wang et al., 2019) laid the groundwork for multi-task evaluation, but were primarily English-centric and biased toward high-resource languages. Frameworks like HELM  (Liang et al., 2022) expanded evaluation to multidimensional axes—accuracy, calibration, robustness, fairness, and efficiency—applied uniformly across 30+ models. Similarly, BIG-Bench (Srivastava et al., 2022) curated over 200 tasks via crowdsourced contributions, highlighting model generality but with limited support for multilingualism or real-world deployment constraints. In parallel, tooling infrastructure also emerged. lm-eval-harness (Gao et al., 2021) provided reproducible few-shot evaluation pipelines, while OpenAI’s Evals (OpenAI Contributors, 2023) offered YAML-based structured evaluation for internal alignment research.
Recent frameworks emphasize modularity, extensibility, and accessibility. For instance, OpenCompass OpenCompass Contributors (2023) integrates distributed inference and real-time leaderboards, though it lacks Indic language support. FreeEval Yu et al. (2024) introduces meta-evaluation and contamination detection, but lacks robust multilingual and Indic language support. DeepEval DeepEval Contributors (2024) supports long-context and tool-use tasks but lacks flexible multilingual and low-resource customization, and its default metrics are often too rigid.
Specialized Capability Benchmarks: To evaluate emerging LLM capabilities, task-specific frameworks and datasets have been proposed. Tool-use and agentic behavior are assessed in ToolBench (Qin et al., 2023) and API-Bank (Li et al., 2023) respectively, which evaluate model interactions with real-world APIs.
Long-context reasoning is explored in InfiniteBench (Zhang et al., 2024) and RULER (Hsieh et al., 2024), which evaluate memory and coherence over 100K+ token sequences. Mathematical reasoning and code generation are evaluated using datasets like GSM8K (Cobbe et al., 2021) and HumanEval (Chen et al., 2021), respectively. While powerful, these benchmarks have limited scope and lack integration into unified evaluation pipelines.
Multilingual and Indic Language Evaluation: Multilingual evaluation has gained prominence through benchmark suites such as XTREME Hu et al. (2020), xP3 Muennighoff et al. (2022), and MEGA Ahuja et al. (2023), covering 40–70 languages. However, these benchmarks often rely on automatic translation, which may not reflect natural language usage or cultural nuances. FLORES Goyal et al. (2022) improved translation evaluation with high-quality parallel corpora for low-resource languages.
For Indian languages, IndicGLUE (Kakwani et al., 2020) and IndicXTREME (Kakwani et al., 2020) pioneered NLU benchmarking across major Indic languages. IndicGenBench (Singh et al., 2024) broadened the scope to generation tasks over 29 Indic languages, but remains a dataset collection rather than a complete evaluation framework, lacking system-level extensibility or plug-and-play usability. Frameworks such as indic-eval (IndicEval Contributors, 2024) wrap existing tools, like LightEval (LightEval, 2024), to support select Indic tasks. However, they offer limited extensibility and lack support for long-context tasks, tool-use evaluation, and custom prompting—features increasingly essential in real-world use cases (Chen et al., 2023).
Eka-Eval is a unified, extensible evaluation framework that integrates the breadth of multilingual and capability-centric benchmarks within a single, production-ready system. It supports a wide spectrum of evaluation settings, including both global and Indic benchmarks spanning 29+ Indian languages, designed for practical deployment and large-scale experimentation.
@@ 3 Capabilities of LLM Evaluation Frameworks
We identify eleven foundational capabilities that are critical for the design and continuous advancement of modern LLM evaluation frameworks. Several of these capabilities are described in earlier evaluation frameworks such as lm-eval-harness (Gao et al., 2021), OpenCompass (OpenCompass Contributors, 2023), FreeEval (Yu et al., 2024), and DeepEval (DeepEval Contributors, 2024). These eleven capabilities are:
(1) Custom Datasets:Support for loading and evaluating user-defined datasets beyond standard benchmark corpora.
(2) Custom Models: Compatibility with a wide range of models, including local checkpoints and API-hosted endpoints.
(3) Custom Prompting: Flexible, template-based prompting paradigms encompassing zero-shot, few-shot, and chain-of-thought configurations.
(4) Long Context: Ability to process and evaluate tasks involving extended input contexts (e.g., exceeding 4,000 tokens).
(5) Tool Use: Evaluating LLMs exhibiting agent-like behavior, including tool use, external API calls, and autonomous multi-step decision-making.
(6) Distributed Inference: Support for parallelized evaluation across multiple processes or compute nodes.
(7) Visual Analysis: Generation of interpretable visualizations, including bar charts, radar plots, and heatmaps, to facilitate comparative analysis.
(8) Multilingual (Indic): Native support for evaluation on benchmarks in Indic languages.
(9) Production Optimization: Implementation of runtime optimizations such as batching, caching, and prompt reuse to enhance evaluation efficiency.
(10) Interactive CLI: Availability of a command-line interface for interactive configuration of datasets, models, prompting strategies, and visualizations.
(11) Quantization: Compatibility with quantized model formats (e.g., 8-bit, 4-bit weights) to minimize memory requirements.
To contextualize the contributions of Eka-Eval, we benchmark it against established alternatives. Table 1 presents a comparative analysis of Eka-Eval and seven widely used LLM evaluation frameworks across the eleven capabilities outlined in this work. Eka-Eval demonstrates robust support across all key capabilities.
@@ 4.1 Design and Implementation
Eka-Eval is architected as a modular, extensible evaluation framework that balances comprehensive benchmark coverage with practical usability. The system is designed around three core principles: modularity for easy extension and customization; accessibility for low-barrier adoption across diverse research environments; and comprehensiveness supports a wide range of capabilities, including underserved areas such as long-context reasoning and tool use.
@@ 4.2 System Architecture
The framework follows a layered architecture consisting of four primary components:
@@ 4.2.1 Evaluation Engine
This component orchestrates all evaluation workflows:
Task Scheduler: Manages task scheduling, prompt formatting, and result aggregation across distributed inference setups. The scheduler implements intelligent work distribution as demonstrated in main_orchestrator() - dynamically assigning evaluation tasks to available workers based on resource constraints and model requirements.
@@
Distributed Coordinator: Coordinates evaluation across multiple GPUs and workers using Python’s multiprocessing library. The coordinator launches multiple worker_process instances to handle independent evaluation tasks, enabling parallel execution across different benchmarks and model configurations.
@@ 4.2.2 Benchmark Registry
Dataset Manager: The BenchmarkRegistry class handles diverse dataset formats and sources, abstracting the complexities of different evaluation protocols. The manager supports datasets from HuggingFace Hub, local files, and custom data formats through standardized interfaces.
@@ 4.2.3 Model Interface Layer
Abstracts access to different local and API-based model backends:
@@
Interactive Selection Interface: Implements get_model_selection_interface() for dynamic model discovery and selection, supporting local model paths and API configurations.
Resource Manager: Ensures efficient memory management through explicit cleanup functions, preventing resource leaks during repeated evaluation runs.
Framework Setup & Config Navigation Result Export Indic Support Extensibility
[…]
@@ 4.2.4 Results Processing System
Metrics Calculator: Computes evaluation metrics using HuggingFace’s evaluate library (e.g., accuracy, BLEU, F1-score, exact match, Pass@1). It also implements robust error handling for edge cases and missing data; for example, when a model returns “The answer is probably B”, regex-based extraction retrieves the label; if that fails, a default score is assigned. Code completions are sandboxed with timeout control to ensure safe execution and error logging.
Visualisations analytics: Provides comparative analysis across multiple models and benchmark configurations by generating visualizations such as bar charts, heatmaps and radar plots (including support for cross-model performance comparisons).
Export Manager: Handles result export in multiple formats including JSON and CSV. The manager maintains evaluation metadata including model parameters, benchmark versions, execution timestamps, and system configurations.

### v4

# Eka-Eval [Uncaptioned image] : An Evaluation Framework for Low-Resource Multilingual Large Language Models

@@ Abstract
The rapid evolution of Large Language Models’ has underscored the need for evaluation frameworks that are globally applicable, flexible, and modular, and that support a wide range of tasks, model types, and linguistic settings. We introduce Eka-Eval, a unified, end-to-end framework that combines a zero-code web interface and an interactive CLI to ensure broad accessibility. It integrates 50+ multilingual benchmarks across nine evaluation categories, supports local and proprietary models, and provides 11 core capabilities through a modular, plug-and-play architecture. Designed for scalable, multilingual evaluation with support for low-resource multilingual languages, Eka-Eval is, to the best of our knowledge, the first suite to offer comprehensive coverage in a single platform. Comparisons against five existing baselines indicate improvements of at least 2x better on key usability measures, with the highest user satisfaction, faster setup times, and consistent benchmark reproducibility.
Website Eka-Eval
Demo bit.ly/Eka-Eval
Code github.com/lingo-iitgn/eka-eval
Eka-Eval [Uncaptioned image] : An Evaluation Framework for Low-Resource Multilingual Large Language Models
@@ 1 Introduction
Large Language Models’ (LLMs) growing capabilities continue to reshape NLP, enabling impressive generalization across diverse tasks, including instruction following, reasoning, summarization, translation, and tool use. With the advent of general-purpose foundation models such as GPT-4 (achiam2023gpt), Claude (anthropic2024claude35sonnet), Gemini (team2023gemini), Deepseek (liu2024deepseek), and Llama-3 (touvron2023llama), the focus of research has increasingly shifted from building task-specific models to systematically evaluating these powerful systems. Evaluation plays a critical role not only in measuring progress but in identifying capabilities, exposing limitations, and informing deployment strategies.
@@
Core Flexibility Advanced Capabilities Usability & Specialization
Framework
Custom
Data
Custom
Models
Custom
Prompt
Quant-
ization
Long
Context
Tool
Use
Dist.
Infer.
Visual
Analysis
Zero-Code
UI
Interact.
CLI
Low-Resource
Multilingual
@@
Table 1: Feature comparison with state-of-the-art frameworks. Takeaway: Eka-Eval seamlessly integrates an accessible dual interface (Zero-Code UI + Interactive CLI) and cutting-edge features (Long Context, Tool Use) with Low-Resource Multilingual benchmarks.
An Alarming Need: Existing evaluation frameworks for LMs, such as HELM (liang2022helm), EleutherAI Harness (gao2021framework), lm-eval-harness (gao2021framework), and OpenCompass (wu2023opencompass), predominantly target high-resource languages and rely on command-line interfaces with limited benchmark coverage (watts2024pariksha). These frameworks create significant barriers for non-technical users who require graphical interfaces and researchers working with low-resource and multilingual settings. The challenge is especially pronounced in linguistically diverse regions globally, where numerous languages serve large populations; yet, existing evaluation frameworks lack comprehensive multilingual support and culturally grounded benchmarks, limiting rigorous assessment of language model capabilities across diverse linguistic settings (pomerenke2025ai; xuan-etal-2025-mmlu; vayani2025all).
Motivation: Existing evaluation frameworks suffer from several limitations such as (1) absence of user-friendly interfaces, requiring extensive command-line expertise, broken-pipelines and manual configuration (gao2021framework; liang2022holistic), (2) restricted flexibility in benchmark selection and task customization (wang2018glue; srivastava2022beyond), (3) insufficient task diversity, with fragmented support for advanced capabilities like long-context reasoning and tool use (qin2023toolllm; zhang2024bench), and (4) limited multilingual coverage, particularly for low-resource Indian languages (hu2020xtreme; ahuja2023mega).
To address this, we propose Eka-Eval, a uniquely scalable and extensible evaluation framework that unifies multilingual benchmarks under flexible task configurations. It enables seamless model integration and multi-platform accessibility, establishing a new benchmark for low-resource multilingual LLM evaluation frameworks.
Contributions: The key contributions are:
@@
  • We propose Eka-Eval, a unified and modular evaluation framework that integrates over 50 multilingual benchmarks across nine major evaluation categories.
  • We introduce a first-ever Zero-Code UI for benchmark evaluation in diverse and low-resource multilingual settings, along with visualisations, leaderboard, and an interactive CLI, making large-scale evaluation accessible to both non-technical and technical users.
  • Through a comprehensive human evaluation involving eleven participants comparing five existing frameworks, showing that Eka-Eval achieves the highest participant ratings across all six features, performing at least 2× better than existing frameworks in key usability and effectiveness.
2 Relevant Work
General-Purpose Frameworks: Early multi-task suites such as GLUE (wang2018glue) and SuperGLUE (wang2019superglue) established foundational standards for evaluation. More comprehensive frameworks, including HELM (liang2022helm), BIG-Bench (srivastava2022beyond), lm-eval-harness (gao2021framework), and OpenAI Evals (openai2023evals) broaden task coverage and introduce extensible evaluation pipelines. yet typically require substantial effort and provide limited multilingual capabilities. Modular ecosystems such as OpenCompass (wu2023opencompass) and FreeEval (yu2024freeeval) offer distributed inference and advanced features, but still lack unified evaluation and zero-code accessibility.
Specialized Capability Benchmarks: Several domain-specific benchmarks target advanced LLM abilities, such as tool use in ToolBench (qin2023toolbench), agentic behavior in API-Bank (li2023api) and InfiniteBench (zhang2024bench), long-context reasoning in RULER (hsieh2024ruler), and mathematical and coding competencies in GSM8K (cobbe2021training) and HumanEval (chen2021evaluating). While powerful, these benchmarks are fragmented across independent repositories, requiring users to mainitain multiple codebases to conduct a thorough evaluation.
Multilingual Evaluation: Multilingual suites such as XTREME (hu2020xtreme) and MEGA (ahuja2023mega) cover 30+ languages but often rely on machine-translated content that lacks cultural grounding. Recent efforts of MCEval (Huang2025MCEvalAD) (13 cultures, 39K dynamic instances), GlotEval (Luo2025GlotEvalAT) (language-agnostic, culturally adaptive prompts), BLEnD (myung2024blend) (16 cultures, 52K native QA pairs), CulturalBench (chiu2024culturalbench) (45 global regions), and ALM-Bench (vayani2025all) (100 languages with cultural images) address this gap. However, these benchmarks still face insufficient data for low-resource languages, cultural biases favor dominant dialects, and evaluation metrics struggle to capture cultural appropriateness.
Eka-Eval mitigates these gaps by unifying 50+ diverse benchmarks within a single modular workflow. It offers dual accessibility (zero-code UI + interactive CLI), advanced capabilities (long-context, tool-use, distributed inference), and multilingual language support for comprehensive evaluation.
@@ 3 Capabilities of LLM Evaluation Frameworks
We identify eleven foundational capabilities critical for modern LLM evaluation frameworks, informed by established systems like lm-eval-harness (gao2021framework), OpenCompass (wu2023opencompass), FreeEval (yu2024freeeval), and DeepEval (deepeval). These capabilities are organized into three categories:
A. Core Flexibility:
(1) Custom Datasets: Support for user-defined evaluation data beyond standard benchmarks.
(2) Custom Models: Compatibility with local checkpoints and API-hosted endpoints.
(3) Custom Prompting: Flexible templates supporting zero-shot, few-shot, and chain-of-thought configurations.
(4) Quantization: Support for memory-efficient quantized formats (8-bit, 4-bit weights).
B. Advanced Capabilities:
(5) Long Context: Processing extended inputs exceeding 4,000 tokens.
(6) Tool Use: Evaluating agent-like behavior including API calls and multi-step reasoning.
(7) Distributed Inference: Parallelized evaluation across multiple compute nodes.
(8) Visual Analysis: Generation of interpretable visualizations (bar charts, radar plots, heatmaps, etc.).
C. Usability & Specialization:
(9) Zero-Code UI: Web-based interface enabling evaluation without programming.
(10) Interactive CLI: Command-line interface for configuring datasets, models, and strategies.
(11) Low-Resource Multilingual: Comprehensive multilingual benchmarks featuring language-specific tokenization.
Table 1 presents a feature-level comparison demonstrating that Eka-Eval is the only framework offering complete coverage across technical, linguistic, and usability dimensions. While lm-eval-harness, OpenCompass, and HELM provide strong technical capabilities, they require substantial expertise and suffer from usability challenges. indic-eval, though tailored to Indic languages, supports only seven benchmarks and lacks broader functionality. Critically, existing frameworks lack zero-code web interfaces, restricting accessibility to users with programming expertise. Eka-Eval uniquely combines comprehensive evaluation capabilities with an intuitive zero-code UI, significantly lowering barriers to rigorous multilingual assessment.
@@ 4.1 Design and Implementation
Eka-Eval is architected as a modular, extensible evaluation framework that balances practical usability with comprehensive benchmark coverage. It provides a zero-code, web-based interface that allows users to run benchmarks, adjust parameters, and visualize results through interactive analytics and leaderboards, while a CLI enables detailed configuration and large-scale evaluations. The framework supports flexible model selection for both local and API-based models and is built around three core principles: modularity for seamless extension and customization; accessibility to support diverse environments, including low-resource language settings; and comprehensiveness to cover a wide range of capabilities and multilingual benchmarks, with particular attention to underserved areas such as long-context reasoning and tool use.
@@ 4.2 System Architecture
The framework has a layered architecture with four main components (illustrated in Figure 1):
@@ 4.2.1 Evaluation Engine
Manages all evaluation workflows:
Task Scheduler: Manages task scheduling, prompt formatting, and result aggregation across distributed inference setups. The scheduler implements intelligent work distribution, as shown in main_orchestrator(), by dynamically assigning evaluation tasks to available workers based on resource constraints and model requirements.
@@
Distributed Coordinator: Coordinates evaluation across multiple GPUs and workers using Python’s multiprocessing library, launching multiple worker_process instances to execute independent evaluation tasks in parallel across benchmarks and model configurations.
@@ 4.2.2 Benchmark Registry
Dataset Manager: The BenchmarkRegistry class handles diverse dataset formats and sources, abstracting the complexities of different evaluation protocols. The manager supports datasets from HuggingFace Hub, local files, and custom API-models through standardized interfaces.
@@ 4.2.3 Model Interface Layer
Access to different local and API-based models:
@@
Interactive Selection Interface: Implements get_model_selection_interface() for dynamic model discovery and selection, supporting local model paths and API configurations(See Appendix C).
Resource Manager: Ensures efficient memory management via cleanup functions, preventing resource leaks during repeated evaluation runs.
Framework Setup & Config Navigation UI Result Export Extensibility Multilingual Support
lm-eval-harness 3.73 $\pm$ 0.72 3.91 $\pm$ 0.66 1.00 $\pm$ 0.00 3.64 $\pm$ 0.70 4.18 $\pm$ 1.00 1.64 $\pm$ 0.50
OpenCompass 2.18 $\pm$ 0.93 2.36 $\pm$ 0.67 3.09 $\pm$ 1.36 3.09 $\pm$ 0.93 2.80 $\pm$ 1.06 2.09 $\pm$ 0.71
HELM 1.91 $\pm$ 0.71 2.64 $\pm$ 1.05 2.36 $\pm$ 1.22 2.55 $\pm$ 1.42 2.09 $\pm$ 0.93 1.36 $\pm$ 0.50
indic-eval 2.55 $\pm$ 0.88 3.18 $\pm$ 0.67 1.00 $\pm$ 0.00 2.55 $\pm$ 1.00 2.55 $\pm$ 0.88 4.55 $\pm$ 0.53
FreeEval 2.55 $\pm$ 1.24 2.45 $\pm$ 0.71 2.73 $\pm$ 1.01 2.64 $\pm$ 1.24 2.36 $\pm$ 0.87 1.45 $\pm$ 0.53
Eka-Eval 4.27 $\pm$ 0.93 4.55 $\pm$ 0.73 4.64 $\pm$ 0.53 4.55 $\pm$ 0.50 4.64 $\pm$ 0.50 4.73 $\pm$ 0.50
Table 2: Average participant ratings of the evaluation frameworks by eleven participants (mean $\pm$ standard deviation; Likert scale: 1–5, 1 = poor and 5 = excellent). Takeaway: Eka-Eval achieves the highest ratings across all functionalities, with participants consistently perceiving it as superior across the six assessed capabilities.
   Framework    Time Taken
   lm-eval-harness    $22\pm 7.58$
   OpenCompass    $36\pm 9.83$
   HELM    $58\pm 5.70$
   IndicEval    $32\pm 4.14$
   FreeEval    $35\pm 7.07$
   Eka-Eval    $\mathbf{11\pm 3.18}$
Table 3: Comparison of time taken for installation and configuration across six frameworks reported across five successful runs on GPU. Takeaway: Eka-Eval records the lowest setup time and exhibits consistently low variance.
@@ 4.2.4 Results Processing System
Metrics Calculator: Employs HuggingFace’s industry-standard evaluate library to compute essential metrics such as accuracy, BLEU, F1-score, exact match, and Pass@1, ensuring consistency with widely adopted benchmarks and standard evaluation practices across diverse LLM tasks.
Visualisations analytics: Provides comparative analysis across multiple models and benchmark configurations by generating visualizations such as bar charts, heatmaps, and radar plots (including support for cross-model comparisons).
Export Manager: Handles result export in multiple formats, including JSON and CSV. The manager maintains evaluation metadata including model parameters, benchmark versions, execution timestamps, and system configurations.
4.2.5 Interface and Deployment Layer
Handles complex evaluation logic with ease of use through a streamlined full-stack user interface (See Appendix §B and Figure 4).
Full-Stack Architecture: A decoupled React frontend and FastAPI backend, served through Nginx, provide a stable and scalable deployment layer.
Real-Time Telemetry: WebSocket-based streaming delivers live inference logs and GPU status directly to the UI, eliminating the need for CLI monitoring.
LLM-Powered Diagnostics: An LLM (LLaMA-3.3 70B (grattafiori2024llama)) interprets evaluation outputs and generates summaries of model performance and failure patterns, revealing additional insights.
Interactive Dashboard: A “Zero-Code” interface with visual comparisons and leaderboards enables users to run and analyze evaluations without technical overhead.